# MTDT: A Multi-Task Deep Learning Digital Twin

### v1

@@ Abstract
Traffic congestion has significant impacts on both the economy and the environment. Measures of Effectiveness (MOEs) have long been the standard for evaluating the level of service and operational efficiency of traffic intersections. However, the scarcity of traditional high-resolution loop detector data (ATSPM) presents challenges in accurately measuring MOEs or capturing the intricate temporospatial characteristics inherent in urban intersection traffic. In response to this challenge, we have introduced the Multi-Task Deep Learning Digital Twin (MTDT) as a solution for multifaceted and precise intersection traffic flow simulation. MTDT enables accurate, fine-grained estimation of loop detector waveform time series for each lane of movement, alongside successful estimation of several MOEs for each lane group associated with a traffic phase concurrently and for all approaches of an arbitrary urban intersection. Compared to existing deep learning methodologies, MTDT distinguishes itself through its adaptability to local temporal and spatial features, such as signal timing plans, intersection topology, driving behaviors, and turning movement counts. While maintaining a straightforward design, our model emphasizes the advantages of multi-task learning in traffic modeling. Although additional complexities could potentially enhance feature learning, we have prioritized simplicity to underscore the benefits of multi-task learning in modeling traffic flow dynamics. By consolidating the learning process across multiple tasks, MTDT demonstrates reduced overfitting, increased efficiency, and enhanced effectiveness through the sharing of representations learned by different tasks. Furthermore, our approach facilitates sequential computation and lends itself to complete parallelization through GPU implementation. This not only streamlines the computational process but also enhances scalability and performance.
@@ I Introduction
Microscopic simulators e.g., VISSIM, SUMO, and AIMSUN fail to generalize to arbitrary topology and infrastructure of traffic urban roadways without painstakingly (re)drawing base maps and (re)setting traffic simulation parameters. Some data sources e.g., GPS, Video, BlueTooth, DSRC etc. are normally sparse with privacy and ownership restrictions [19]. Advanced sensing technologies such as induction loop detectors (ATSPM) make high-resolution loop detector traffic data widely available, although it fails to provide enough information for MOEs estimation by itself. The computation of MOEs is specific to several tempo-spacial factors such as intersection characteristics(e.g., topology, lane configuration), traffic pattern (e.g., inflow traffic and turning-movement counts, signal timings, and traffic conditions (e.g., driving behaviors).
@@
Existing analytical methods of estimating MOEs from ATSPM either use vehicle conservation input-output algorithms based on counting vehicle flow over advance and stop-bar detectors [18] or Lighthill’s shockwave algorithms based on identifying breakpoints in the detector actuation waveforms [22]. Although, both of them are sensitive to the placement of detectors and limited to the localized training datasets. Machine learning methods use either GPS trace data [5] that are too coarse in recording dissipation trends, or ATSPM for cycle-wise MOEs which is not based on real-world traffic conditions due to the random vehicle flow and road networks without explicit left-turn buffers. Recently, some deep learning methods have employed geometric deep learning [20] and graph convolution networks [7] to extract cycle-wise MOEs, although these approaches still fail to generalize the local traffic conditions and road geometries, also their reconstructed MOEs have as low resolution as approach level.
@@
We propose a Multi-Task Deep Learning-based (MTDT) digital twin of intersections to address the aforementioned shortcomings. This model is trained on real-world data obtained from high-resolution induction loop detectors (ATSPM), capturing varying traffic patterns encompassing a wide range of configurable parameters, thus generating potentially viable counterfactual traffic scenarios. MTDT builds upon the previous work by Yousefzadeh et al. [23], extending it to create an accurate and reliable multi-purpose digital twin. An overview of our model is depicted in Figure 1, where orange-colored objects represent the inputs of the model, including stop-bar loop detector waveforms and signal timing plans (additinally are present driving behavior parameters and turning-movement counts features), while blue-colored objects represent the outputs of the model, including exit and inflow waveforms at each traffic lane, maximum queue length at each traffic phase, and travel time distribution at each traffic phase. Further details regarding the definition and computation of these MOEs are provided in Section IV. Traffic Data Generation.
@@
  • MTDT is a multi-tasking digital twin engineered to comprehensively capture the nuanced facets of traffic dynamics as vehicles approach and depart from any signalized intersection. Its outputs encompass fine-grained and lane-wise estimations of two primary tasks: exit and inflow waveform time series, along with accurate and phase-wise estimations of two secondary tasks: maximum queue length waveforms and travel time distributions associated with every phase of movement. The model’s adaptability and responsiveness to the temporospatial intricacies of traffic dynamics are evident in its utilization of input parameters such as relevant driving behavior metrics, signal timing plans, turning movement counts, and signal timing plan parameters, among others.
  • Using our digital twins, it becomes feasible to accurately and reliably predict the impact of loop detector waveforms, signal timing plans, driving behavior parameters, and turning movement counts on lane-wise platoons and related MOEs. Importantly, this predictive capability is achieved at a significantly faster pace compared to microscopic simulators, thanks to the utilization of GPU-based processors. The methodologies proposed in our digital twins incur $O(1)$ sequential computation and can be entirely parallelized using cost-effective GPU computation.
  • The findings of this study hold significance for making data-driven decisions regarding changes in traffic signal timing, formulating traffic strategies, planning transportation systems, and redesigning the infrastructure of roadways and intersections. Our digital twins offer valuable insights and tools for traffic control design and safety management, particularly in the context of smart intersections operating within highly dynamic environments.
We designed several experiments, whose findings aim to showcase the feasibility and efficacy of our multi-tasking approach in designing a deep learning-based and multifaceted digital twin for intersections. We introduce two variants of our digital twin and compare our model with them and with our previously introduced TQAM [15] digital twin. The results of the experiments, underscore the versatility of our proposed digital twin, demonstrating its ability to not only generalize estimations across various intersection typologies and characteristics, but also can successfully simulating ATSPM time series waveforms at lane level of resolution, and accurately estimating several Measures of Effectiveness (MOEs) in both time series and distribution forms at phase (lane-group) of resolution for every approach (movement direction) within an urban intersection.
@@
The rest of the paper is organized as follows: Section VI provides an overview of related work, discussing various techniques employed for traffic state estimation. Section III outlines the architecture of our proposed digital twin. Finally, in Section V, we assess and compare the performance of our model through a series of experiments involving various observable and unobservable quantities.
@@ II Problem Definition
Definition 1. Simulation Record: This term refers to a record of a simulation run represented as a tuple of intersection ID, signal waveform $sig$, turn-movement counts vector $tmc$, driver behavior vector $drv$, lane-wise stop-bar loop detector waveforms $stp$, exit loop detector waveforms $ext$, and inflow loop detector waveforms $inf$, and phase-wise queue length waveforms $ql$, and travel time histograms $tt$.
@@
Notation Description Aggr Size Type $j$ Intersection - 1x9 String $stp$ Waveform at Stop-bar detector 5 sec 48x80 Integer, 0-8 $ext$ Waveform at the exit of the intersection 5 sec 16x80 Integer, 0-8 $inf$ Waveform upstream the intersection 5 sec 12x80 Integer, 0-8 $ql$ Queue length time series 5 sec 8x80 Integer, 0-1200 $tt$ Travel time histogram 5 sec 8x200 Integer, 0-700 $sig$ Signal timing state information 5 sec 8x80 Binary $drv$ Driving behavior parameters 2400 sec 1x9 Float, 0-30 $tmc$ Turning-movement counts ratio 2400 sec 35x35 Float, 0-1 $w$ Size of the prediction window - 1 Integer, 0-150 $s$ Single simulation recorded by SUMO 5 sec - Multiple $A$ Common lane-connectivity for $M_{ext}$ module - 2x22 Binary $A^{\prime}$ Common lane-connectivity for $M_{inf}$ module - 2x72 Binary $v$ multivariate time series for queue length estimation 5 sec 1x7x80 - $v^{\prime}$ multivariate time series for travel time estimation 5 sec 1x7x80 - $M_{ext}$ GNN module for exit waveforms estimation - - torch.nn.Module $M_{inf}$ GNN module for Inflow waveforms estimation - - torch.nn.Module $M_{ql}$ CNN module for queue length estimation - - torch.nn.Module $M_{tt}$ CNN module for travel time estimation - - torch.nn.Module
@@ III Proposed Models
In this section, we present our innovative deep-learning solution for a versatile and multi-purpose traffic intersection digital twin. Our Multi-task Deep Learning Digital Twin (MTDT) is designed to simultaneously grasp various facets of intersection-level traffic flow dynamics. MTDT is jointly optimized to learn four sets of joint tasks. The general architecture of our digital twin is composed of two primary modules that yield base learning representations, and two secondary modules that use the representations learned by primary modules to learn their own tasks and improve the learning process of primary tasks as well by the sharing of their representations. For a comprehensive understanding of MTDT’s architecture, refer to the overview depicted in Figure 2.
Figure 2: Overview of the architecture of MTDT. The architecture of our proposed Multi-componential digital twin framework shows two types of modules are interconnection within the pipeline. GATs modules (highlighted in light orange) primarily execute two tasks, while CNNs modules (depicted in dark blue) are tailored for secondary tasks. During inference mode, red-colored arrows activate to seamlessly integrate the aggregated output of primary modules into the input of secondary modules.
The MTDT pipeline comprises two distinct types of modules: Graph Attention Networks (GATs) [17] modules, including $M_{ext}$ and $M_{inf}$, dedicated to the primary task of estimating loop detector waveforms, and Convolutional Neural Networks (CNNs) [8] modules, including $M_{ql}$ and $M_{tt}$, geared towards secondary tasks such as estimating Measures of Effectiveness (MOEs) like maximum queue length and travel time distribution. More specifically, GATs modules, namely $M_{ext}$ and $M_{inf}$, excel at precise, lane-specific estimations of waveform time series for vehicle platoons approaching and departing the target intersection. They meticulously consider instantaneous temporospatial factors like turning movement counts, driving behavior parameters, and signal timing plans. Subsequently, CNNs modules, such as $M_{ql}$ and $M_{tt}$, utilize the outputs from primary modules alongside instantaneous temporospatial characteristics of the target intersection to estimate MOEs like maximum queue length and travel time distribution. This design ensures adaptability without restrictive assumptions, making the digital twin easily instantiated for various secondary tasks related to traffic dynamics.
@@
Detailed information regarding the architecture and learning process within each module is elaborated upon in the subsequent subsections:
@@ III-A Exit Module
Where $N_{i}$ is the connected (neighboring node) lanes of each lane $i$ in the intersection’s graph representation. A fully connected layer activated by a ReLU function is applied on the output of the GATConv layer to impute missing (unobserved) time series waveforms for exit loop detectors $ext$ within reconstructed input feature matrix $\hat{X}\in\mathbb{R}^{N\times w}$.
@@ III-B Inflow Module
Inflow subnetwork module $M_{inf}$ is an attentional graph neural network (GAT) with an imputation mechanism, as a simplified version of multi-layer graph neural network [24] solution introduced in [23]. This modules is responsible to perform the primary task of inflow waveform time series $inf$ estimation. The difference between $M_{ext}$ and $M_{inf}$ is that the same GATs module architecture is adopted to perform a different primary task. Here, GATConv implicitly specifies and integrates the attention coefficients denoting the importance $stp$ waveform time series to each $inf$ waveform using a different graph representation of an intersection that connects each of 12 inflow loop detectors’ $ext$ waveform time series (for 3 inflow lanes for each approach) to the total number of 48 stop-bar loop detectors’ $stp$ waveform time series as in a general topology of an arbitrary up to 4-way intersection. In this setting, this GATs module can impute missing (unobserved) time series waveforms for exit loop detectors $inf$ within reconstructed input feature matrix $\hat{X}\in\mathbb{R}^{N^{\prime}\times w}$.
@@ III-C Queue Length Module
Queue Length subnetwork module $M_{ql}$ is a multi-output 1D convolutional neural network (CNN) to perform the secondary task of maximum queue length waveform time series $ql$ estimation. Here, the input data is fed in the form of Multivariate Time series $v_{s,j}$ of size $(8,7,80)$ each time step for each of 8 traffic phase, pair 7 variables including exit waveform time series $ext$ estimated by $M_{ext}$ with $stp$, driving behavior parameters $drv$, and signal timing plans $sig$ to capture appropriate and instant tempospatial characteristics of urban traffic intersections. In this setting, for simplicity reasons, only the three most effective driving behavior i.e., accel, lcCooperative, minGap in traffic dynamic of intersection as shown in [23]. The architecture of the CNNs module consists of 3 Conv1d layers each followed by a MaxPool1d layer and flattening layer before being fed into a fully connected layer activated by a ReLU non-linearity to simultaneously generate a set of 8 outputs as an estimation to maximum queue length waveform time series for every phase of target standard NEMA eight-phasing intersection.

### v2

@@ Abstract
Traffic congestion has significant impacts on both the economy and the environment. Measures of Effectiveness (MOEs) have long been the standard for evaluating traffic intersections’ level of service and operational efficiency. However, the scarcity of traditional high-resolution loop detector data (ATSPM) presents challenges in accurately measuring MOEs or capturing the intricate spatiotemporal characteristics inherent in urban intersection traffic. To address this challenge, we present a comprehensive intersection traffic flow simulation that utilizes a multi-task learning paradigm. This approach combines graph convolutions for primary estimating lane-wise exit and inflow with time series convolutions for secondary assessing multi-directional queue lengths and travel time distribution through any arbitrary urban traffic intersection. Compared to existing deep learning methodologies, the proposed Multi-Task Deep Learning Digital Twin (MTDT) distinguishes itself through its adaptability to local temporal and spatial features, such as signal timing plans, intersection topology, driving behaviors, and turning movement counts. We also show the benefit of multi-task learning in the effectiveness of individual traffic simulation tasks. Furthermore, our approach facilitates sequential computation and provides complete parallelization through GPU implementation. This not only streamlines the computational process but also enhances scalability and performance.
@@ I Introduction
Microscopic simulators e.g., VISSIM, SUMO, and AIMSUN fail to generalize to arbitrary topology and infrastructure of traffic urban roadways without painstakingly (re)drawing base maps and (re)setting traffic simulation parameters. Some data sources e.g., GPS, Video, BlueTooth, DSRC, etc. are normally sparse with privacy and ownership restrictions [19]. Advanced sensing technologies such as induction loop detectors (ATSPM), make high-resolution loop detector traffic data widely available, although they fail to provide enough information for MOEs estimation by themselves. The computation of MOEs is specific to several tempo-spacial factors such as intersection characteristics(e.g., topology, lane configuration), traffic pattern (e.g., inflow traffic and turning-movement counts, signal timings, and traffic conditions (e.g., driving behaviors).
@@
Existing analytical methods of estimating MOEs from ATSPM either use vehicle conservation input-output algorithms based on counting vehicle flow over advance and stop-bar detectors [18] or Lighthill’s shockwave-based methods that identify breakpoints in the detector actuation waveforms [22]. Although, both of them are sensitive to the placement of detectors and limited to the localized training datasets. Machine learning methods use either GPS trace data [5] that are too coarse in recording dissipation trends, or ATSPM for cycle-wise MOEs which is not based on real-world traffic conditions due to the random vehicle flow and road networks without explicit left-turn buffers. Recently, some deep learning methods have employed geometric deep learning [20] and graph convolution networks [7] to extract cycle-wise MOEs, although these approaches still fail to generalize the local traffic conditions and road geometries, also their reconstructed MOEs have as low resolution as approach level.
@@
We propose a Multi-Task Deep Learning-based (MTDT) digital twin of intersections to address the aforementioned shortcomings. This model is trained on real-world data obtained from high-resolution induction loop detectors (ATSPM), capturing varying traffic patterns encompassing a wide range of configurable parameters, thus generating potentially viable counterfactual traffic scenarios. MTDT builds upon the previous work by Yousefzadeh et al. [23], extending it to create an accurate and reliable multi-purpose digital twin. The main difference between them is that [23] introduces two individual deep learning-based digital twins that utilize graph neural networks to make a lane-wise estimation of either exit or inflow waveforms at any direction of an arbitrary intersection. We use a multi-task learning paradigm to combine graph convolution for learning primary lane-wise traffic waveforms with time series convolution for learning secondary multi-directional MOEs based on primary results for an arbitrary urban traffic intersection, all at the same time. While maintaining a straightforward design, our model emphasizes the advantages of multi-task learning in traffic modeling. Although additional complexities could potentially enhance feature learning, we have prioritized simplicity to underscore the benefits of multi-task learning in modeling traffic flow dynamics. By consolidating the learning process across multiple tasks, MTDT demonstrates reduced overfitting, increased efficiency, and enhanced effectiveness through the sharing of representations learned by different tasks. An overview of our model is depicted in Figure 1, where orange-colored objects represent the inputs of the model, including stop-bar loop detector waveforms and signal timing plans (additionally are present driving behavior parameters and turning-movement counts features), while blue-colored objects represent the outputs of the model, including exit and inflow waveforms at each traffic lane, maximum queue length at each traffic phase, and travel time distribution at each traffic phase. Further details regarding the definition and computation of these MOEs are provided in Section IV. Traffic Data Generation.
@@
  • MTDT jointly estimates fine-grained and lane-wise exit and inflow waveform time series, along with accurate travel time distribution and maximum queue lengths associated with every phase of the movement. The model’s adaptability and responsiveness to the spatiotemporal intricacies of traffic dynamics are evident in its utilization of input parameters such as relevant driving behavior metrics, signal timing plans, turning movement counts, and signal timing plan parameters, among others.
  • Using our digital twins, it becomes feasible to accurately and reliably predict the impact of loop detector waveforms, signal timing plans, driving behavior parameters, and turning movement counts on lane-wise platoons and related MOEs. Importantly, this predictive capability is computationally faster than microscopic simulators, thanks to the utilization of GPU-based processors. The methodologies proposed in our digital twins incur $O(1)$ sequential computation and can be entirely parallelized using cost-effective GPU computation.
  • The findings of this study hold significance to extend to corridor and network level digital twins, and make benefits towards data-driven traffic control design and safety management, particularly in the context of smart intersections operating within highly dynamic environments.
We designed several experiments, whose findings aim to showcase the feasibility and efficacy of our multi-tasking approach in designing a deep learning-based and multifaceted digital twin for intersections. We introduce two variants of our digital twin and compare our model with them and with our previously introduced TQAM [15] digital twin. The results of the experiments, underscore the versatility of our proposed digital twin, demonstrating its ability to not only generalize estimations across various intersection typologies and characteristics, but also successfully simulate ATSPM waveforms at lane level, and accurately estimate several Measures of Effectiveness (MOEs) in both time series and distribution forms at phase (lane-group) level for every approach (movement direction) within an urban intersection.
@@
The rest of the paper is organized as follows: Section VI provides an overview of related work, discussing various techniques employed for traffic state estimation. Section III outlines the architecture of our proposed digital twin. Finally, in Section V, we assess and compare the performance of our model through a series of experiments.
@@ II Problem Definition
Definition 1. Simulation Record: This term refers to a record of a simulation run represented as a tuple of intersection ID, signal waveform $sig$, turn-movement counts vector $tmc$, driver behavior vector $drv$, lane-wise stop-bar loop detector waveforms $stp$, exit loop detector waveforms $ext$, and inflow loop detector waveforms $inf$, and multi-directional queue length waveforms $ql$, and travel time histograms $tt$.
@@
Notation Description Aggr Size Type $j$ Intersection - 1x9 String $stp$ Waveform at Stop-bar detector 5 sec 48x80 Integer, 0-8 $ext$ Waveform at the exit of the intersection 5 sec 16x80 Integer, 0-8 $inf$ Waveform upstream the intersection 5 sec 12x80 Integer, 0-8 $ql$ Queue length time series 5 sec 8x80 Integer, 0-1200 $tt$ Travel time histogram 5 sec 8x200 Integer, 0-700 $sig$ Signal timing state information 5 sec 8x80 Binary $drv$ Driving behavior parameters 2400 sec 1x9 Float, 0-30 $tmc$ Turning-movement counts ratio 2400 sec 35x35 Float, 0-1 $w$ Size of the prediction window - 1 Integer, 0-150 $s$ Single simulation recorded by SUMO 5 sec - Multiple $A$ Common lane-connectivity for $M_{ext}$ module - 2x22 Binary $A^{\prime}$ Common lane-connectivity for $M_{inf}$ module - 2x72 Binary $v$ multivariate time series for queue length estimation 5 sec 1x7x80 - $v^{\prime}$ multivariate time series for travel time estimation 5 sec 1x7x80 - $M_{ext}$ GAT module for exit waveforms estimation - - torch.nn.Module $M_{inf}$ GAT module for Inflow waveforms estimation - - torch.nn.Module $M_{ql}$ CNN module for queue length estimation - - torch.nn.Module $M_{tt}$ CNN module for travel time estimation - - torch.nn.Module
@@ III Proposed Models
The Multi-task Deep Learning Digital Twin (MTDT) is designed to grasp various facets of intersection-level traffic flow dynamics simultaneously. MTDT is jointly optimized to learn four sets of joint tasks. The general architecture of our digital twin is composed of two primary modules that yield base learning representations, and two secondary modules that use the representations learned by primary modules to learn their own tasks and improve the learning process of primary tasks as well by the sharing of their representations. For a comprehensive understanding of MTDT’s architecture, refer to the overview depicted in Figure 2. Here, as described in Table 1, $M_{ext}$ and $M_{inf}$ represent GAT modules that execute exit waveforms and inflow waveforms estimation respectively; while $M_{ql}$ and $M_{tt}$ represents CNN modules that execute queue length time series and travel time distribution estimation respectively. The output of primary tasks of size $16\times 80$ for the exit task and $12\times 80$ for the inflow task are aggregated to 8 lane groups (phase of movement) to the size $8\times 80$ before being fed as input to the secondary tasks.
Figure 2: Overview of the architecture of MTDT. The architecture of the proposed Multi-componential digital twin framework contains two types of modules. GAT modules (light orange colored) execute primary tasks, which aggregated outputs are used as input to secondary tasks executed by CNN modules (dark blue colored). Red-colored arrows are activated only in inference mode.
The MTDT pipeline comprises two distinct types of modules: Graph Attention Networks (GATs) [17] modules, including $M_{ext}$ and $M_{inf}$, dedicated to the primary task of estimating loop detector waveforms, and Convolutional Neural Networks (CNNs) [8] modules, including $M_{ql}$ and $M_{tt}$, geared towards secondary tasks such as estimating Measures of Effectiveness (MOEs) like maximum queue length and travel time distribution. More specifically, GAT modules, namely $M_{ext}$ and $M_{inf}$, excel at precise, lane-specific estimations of waveform time series for vehicle platoons approaching and departing the target intersection. They meticulously consider instantaneous spatiotemporal factors like turning movement counts, driving behavior parameters, and signal timing plans. Subsequently, CNN modules, such as $M_{ql}$ and $M_{tt}$, utilize the outputs from primary modules alongside instantaneous spatiotemporal characteristics of the target intersection to estimate MOEs like maximum queue length and travel time distribution. This design ensures adaptability without restrictive assumptions, making the digital twin easily instantiated for various secondary tasks related to traffic dynamics.
@@ III-A Exit Module
Where $N_{i}$ is the connected (neighboring node) lanes of each lane $i$ in the intersection’s graph representation. A fully-connected layer activated by a ReLU function is applied on the output of the GATConv layer to impute missing (unobserved) time series waveforms for exit loop detectors $ext$ within reconstructed input feature matrix $\hat{X}\in\mathbb{R}^{N\times w}$.
@@ III-B Inflow Module
Inflow subnetwork module $M_{inf}$ is an attentional graph neural network (GAT) with an imputation mechanism, as a simplified version of multi-layer graph neural network [24] solution introduced in [23]. This module is responsible for performing the primary task of inflow waveform time series $inf$ estimation. The difference between $M_{ext}$ and $M_{inf}$ is that the same GATs module architecture is adopted to perform a different primary task. Here, GATConv implicitly specifies and integrates the attention coefficients denoting the importance $stp$ waveform time series to each $inf$ waveform using a different graph representation of an intersection that connects each of 12 inflow loop detectors’ $ext$ waveform time series (for 3 inflow lanes for each approach) to the total number of 48 stop-bar loop detectors’ $stp$ waveform time series as in a general topology of an arbitrary up to 4-way intersection. In this setting, this GATs module can impute missing (unobserved) time series waveforms for exit loop detectors $inf$ within reconstructed input feature matrix $\hat{X}\in\mathbb{R}^{N^{\prime}\times w}$.
@@ III-C Queue Length Module
Queue Length subnetwork module $M_{ql}$ is a multi-output 1D convolutional neural network (CNN) to perform the secondary task of maximum queue length waveform time series $ql$ estimation. Here, the input data is fed in the form of Multivariate Time series $v_{s,j}$ of size $(8,7,80)$ each time step for each of 8 traffic phase, pair 7 variables including exit waveform time series $ext$ estimated by $M_{ext}$ with $stp$, driving behavior parameters $drv$, and signal timing plans $sig$ to capture appropriate and instant spatiotemporal characteristics of urban traffic intersections. In this setting, for simplicity reasons, only the three most effective driving behaviors i.e., accel, lcCooperative, minGap in traffic dynamic of intersection, based on SHAP analysis of the Average impact of features on the magnitude of estimation in [23], are included. The architecture of the CNNs module consists of 3 Conv1d layers each followed by a MaxPool1d layer and flattening layer before being fed into a fully connected layer activated by a ReLU non-linearity to simultaneously generate a set of 8 outputs as an estimation to maximum queue length waveform time series for every phase of target standard NEMA eight-phasing intersection.